Grade 12Mechanics

What is the SI unit of gas constant? And what is its CGS unit?Write in different lines.

Kaviya
8 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
  1. Value of  R in M.K.S. System

According to Avogadro’s law, the volume of one mole of a gas at NTP is 22.4 litre. Standard temperature & pressure at N.T.P. are 00C & one atmosphere

P=1 atm.

V=22.4 l

n= 1 mole.

T=  2730K

R=0.0821 litre atmosphere/degree/mole.

 

  1. Value of R. in C.G.S. System

If P is expressed in dyne/Cm2 & V in Cm3

P=1 atm = 76 cm. of Hg = 76×13.6×981 dyne/cm2

V= 22.4 litre = 22400 ml = 22400 cm3

n= 1 mole

T= 0+273 = 2730K

= 8.3 ×107 erg/degree/mole

= 8.3 Joule/degree/mole

Because one calorie is equivalent to 4.08×107erg. So

R =[ 8.3×10erg/degree/mole ×  1calorie] /( 4.08×107 erg )

= 2 calorie/degree/mole